Alfred wants to invest $4,000 at 6% simple interest rate for 5 years. How much interest will he receive?
Answer by checkley77(12844)

4,000(1+.06)^5=4,000(1.06)^5=4,000*1.338=$5,3552.90-$4,000=$1,352.90 is the interest.